OpenStudy (anonymous):

A crowbar is used as a lever. The effort force of 40 newtons moves 3 meters. The resistance force of 54 newtons moves 2 meters. What is the efficiency of the lever? 100% because energy must be conserved 90% 25% 18%

OpenStudy (anonymous):

e = work output / work input work = force * distance e = (54)(2) / (40)(3) = 108/120 = 0.9 = 90%